3(x+4)=3(x-2) please help and response quickly thanks;)
Answer by stanbon(75887)

3(x+4)=3(x-2)
3x+12 = 3x-6
Subtract 3x from both sides:
12 = -6
----
That is a contradiction telling you
the original statement of the problem
is a contradiction. It is also telling
you there is no solution.
============
Cheers,
Stan H.